OpenStudy (anonymous):

if a^2+b^2=8and (a-b)^2=4,what is the value of ab

OpenStudy (anonymous):

come to twiddla http://www.twiddla.com/584457

myininaya (myininaya):

(a-b)^2=a^2-2ab+b^2=a^2+b^2-2ab=8-2ab=4 8-2ab=4 -2ab=-4 ab=2

OpenStudy (anonymous):

at least give a meda iman :)

myininaya (myininaya):

medal party!

OpenStudy (anonymous):

yup...lolx

OpenStudy (anonymous):

you guys are too fast! lolol

myininaya (myininaya):

joe i need your help if you can

OpenStudy (anonymous):

whats up?